The length of a rectangle 6 cm more than the width. If perimeter 48 cm,what are the dimensions?
Answer by bonster(299)

P=2L+2W
L=6 cm more than the width
then width =w and length=w+6
p=2L+2W
48=2(w+6)+2(w)
48=2w+12+2w
48=4w+12
36=4w
w=9
since width=w, the width is 9 cm
since length =(w+6), the length is 15 cm.